OpenStudy (anonymous):

In the problem (-2xy^3)(-4x^2y^2) , Jose got -6x^2y^6 . Correct his 3 mistakes.

Parth (parthkohli):

First: \(-4 \times -2 \ne -6\)

Parth (parthkohli):

Second: \(x \times x^2 \ne x^2\)

Parth (parthkohli):

Third: \(y^3 \times y^2 \ne y^6\)

Parth (parthkohli):

I leave upon you to correct those mistakes.

OpenStudy (anonymous):

what is ne?

Parth (parthkohli):

"Not equal to"

OpenStudy (anonymous):

ok